a coffee barista is mixing two types of coffee beans;el-cheapo coffee beans ( $2 per pound) and premium star-buckaroo coffee beans ($8 per pound) she has 12 pounds of star-buckaroo coffee beans and wishes to mix in a quantity of el-cheapo coffee beans to get a blend that can be sold for $6 per pound.
How many pounds of el-ceapo coffee should be used to create the mixture?

Answer by mananth(16946)

------------------------ price ---------------- quantity
Buckaroo coffee 8 ----------------12 pounds
Elcheapoo coffee 2----------------x pounds
Mixture 6 ---------------- 12 + x pounds
The sum of idividual costs of the two types of coffee = cost of mixture
8*12+2x =6(12+x)
96+2x=72+6x
2x-6x=72-96
-4x=-24
/-4
x=6 pounds of Elcheapoo coffee

m.ananth@hotmail.ca